# Resilience Inference for Supply Chains with Hypergraph Neural Network

### v1

@@ Introduction
We represent the supply chain as a tripartite hypergraph $\mathcal{H}=(\mathcal{C},\mathcal{P},\mathcal{E})$, where $\mathcal{C}$ denotes the set of firm nodes, $\mathcal{P}$ is the set of product nodes, and $\mathcal{E}\subseteq\mathcal{C}\times\mathcal{P}\times\mathcal{C}$ represents the set of hyperedges. Each hyperedge $(c_{\text{up}},p,c_{\text{down}})\in\mathcal{E}$ indicates that the downstream firm $c_{\text{down}}$ procures product $p$ from the upstream firm $c_{\text{up}}$. Additionally, each firm node $c\in\mathcal{C}$ is associated with a time-dependent state vector $\smash{\mathbf{x}_{c}^{(t)}}\in\mathbb{R}^{d}$, capturing its operational status at time $t$. The dimensions of this vector may include key performance indicators such as inventory levels, production rates, and order volumes. In this work, we use inventory vectors to define node states: $\mathbf{x}_{c}^{(t)}=\bigl(I_{c,p}^{(t)}\bigr)_{p\in\mathcal{P}}\in\mathbb{R}^{|\mathcal{P}|}$, where $I_{c,p}^{(t)}$ denotes the inventory level of product $p$ held by firm $c$ at time $t$. Stacking these vectors across all firms forms the supply chain state matrix $\mathbf{X}_{t}\in\mathbb{R}^{|\mathcal{C}|\times|\mathcal{P}|}$.
@@ Acknowledgments
This work was partially supported by the grants of Jilin Provincial International Cooperation Key Laboratory for Super Smart City and Jilin Provincial Key Laboratory of Intelligent Policing.
Appendix A Appendix
This appendix describes the discrete-time simulator used to generate resilience labels for firm–product networks. The simulation follows Forrester’s stock–flow feedback framework (forrester1997industrial) and incorporates Sterman’s exponential-smoothing demand model with bullwhip amplification effects (sterman2002system). Updates proceed at discrete time steps $t=0,1,2,\dots$, using a unit interval.
The simulator first generates inventory trajectories over time under dynamic supply–demand interactions. Final convergence behavior under varying initial conditions is then used to derive resilience labels.
Notation
Table 1 summarizes the variables used. Subscripts $c$ and $p$ denote a firm and its product, respectively; $r$ indicates raw material components.
Symbol Definition
$\beta$ Exponential-smoothing coefficient (default: 0.2)
$\theta_{c,p}$ Inventory coverage period (default: 1)
$\tau_{c,p}$ Gap-closing horizon in time steps (default: 1)
$\alpha_{c,p,r}$ BOM coefficient (units of $r$ per unit $p$ at $c$)
$\hat{D}_{c,p}(t)$ Forecast demand for $(c,p)$ at time $t$
$D^{\text{ext}}_{c,p}(t)$ External market demand
$O_{c,p}(t)$ Order rate placed for product $p$
$I_{c,p}(t)$ Inventory of product $p$ at firm $c$
$I^{*}_{c,p}(t)$ Target inventory level
$U_{c,p}(t)$ Inventory gap
$Q^{\text{req}}_{c,p}(t)$ Requested production rate
$Q^{\max}_{c,p}(t)$ Feasible production upper bound
$Q^{\text{plan}}_{c,p}(t)$ Final planned production rate
$C_{c,r}(t)$ Consumption rate of raw material $r$
Table 1: Symbols used in the inventory–production simulation.
Simulation Procedure
We now describe the seven-step simulation logic.
1. Aggregated demand.
Each firm faces demand from both the external market and downstream customers. We fix the exogenous demand for the most downstream firms at 100 to isolate internal inventory dynamics:
$D^{\text{agg}}_{c,p}(t)=D^{\text{ext}}_{c,p}(t)+\sum_{d\in\mathrm{down}(c)}O_{d,p}(t)$ (1)
where $\mathrm{down}(c)$ is the set of firms directly downstream of $c$.
2. Demand forecast.
Demand is forecast using exponential smoothing with coefficient $\beta=0.2$, balancing responsiveness and stability:
$\hat{D}_{c,p}(t)=\beta\,D^{\text{agg}}_{c,p}(t)+(1-\beta)\,\hat{D}_{c,p}(t-1)$ (2)
3. Target inventory and gap.
Each firm seeks to maintain inventory for $\theta_{c,p}$ forecast periods. The target and gap are:
$I^{*}_{c,p}(t)$ $=\theta_{c,p}\cdot\hat{D}_{c,p}(t),$ (3)
$U_{c,p}(t)$ $=\max\{0,\,I^{*}_{c,p}(t)-I_{c,p}(t)\}$
In our experiments, we use a constant $\theta_{c,p}=1$ for all firm–product pairs, ensuring that firms aim to hold inventory equivalent to one forecast period.
4. Requested production.
To close the inventory gap uniformly over $\tau_{c,p}$ future steps:
$Q^{\text{req}}_{c,p}(t)=\frac{U_{c,p}(t)}{\tau_{c,p}}$ (4)
We fix $\tau_{c,p}=1$ in all simulations, reflecting an aggressive replenishment policy that seeks to eliminate gaps in a single step.
5. Raw-material feasibility.
Requested production may not be fully realizable due to raw material constraints. Each product $p$ requires a set of raw inputs $\mathcal{R}(c,p)$ according to a bill-of-materials (BOM), with $\alpha_{c,p,r}$ denoting the required units of raw material $r$ per unit of product $p$ at firm $c$. All BOM coefficients are drawn from the SupplySim simulator (chang2025learning).
At each time step, the feasible production rate is determined by the most limiting input:
$Q^{\max}_{c,p}(t)=\min_{r\in\mathcal{R}(c,p)}\frac{I_{c,r}(t)}{\alpha_{c,p,r}}$ (5)
The final planned production rate is then set as the minimum between the requested and feasible rates:
$Q^{\text{plan}}_{c,p}(t)=\min\{Q^{\text{req}}_{c,p}(t),\,Q^{\max}_{c,p}(t)\}$ (6)
This ensures that planned production remains within material availability while still addressing the inventory gap.
6. Raw-material dynamics.
Once the planned production rate $Q^{\text{plan}}_{c,p}(t)$ is determined, it induces corresponding demand and consumption for each raw material $r\in\mathcal{R}(c,p)$. The raw material forecast, target inventory, gap, order, and consumption are computed analogously to finished products:
$\hat{D}_{c,r}(t)$ $=\alpha_{c,p,r}\cdot\hat{D}_{c,p}(t)$ (7)
$I^{*}_{c,r}(t)$ $=\theta_{c,r}\cdot\hat{D}_{c,r}(t)$ (8)
$U_{c,r}(t)$ $=\max\{0,\,I^{*}_{c,r}(t)-I_{c,r}(t)\}$ (9)
$O_{c,r}(t)$ $=\frac{U_{c,r}(t)}{\tau_{c,r}}$ (10)
$C_{c,r}(t)$ $=\alpha_{c,p,r}\cdot Q^{\text{plan}}_{c,p}(t)$ (11)
Actual upstream inflows of raw materials are constrained by network-level availability and routing. Each upstream firm evenly distributes its available outflow to all connected downstream receivers. The realized inflow $\mathcal{I}_{c,r}(t)$ is then the minimum between the incoming supply and the placed order:
$\mathcal{I}_{c,r}(t)=\min\left\{\sum_{u\in\mathcal{U}(c,r)}\frac{\mathrm{out}_{u,r}(t)}{N_{u,r}},\;O_{c,r}(t)\right\}$ (12)
where $\mathcal{U}(c,r)$ is the set of upstream suppliers of raw material $r$ to firm $c$, and $N_{u,r}$ is the number of downstream firms they serve. For top-layer firms with no upstream suppliers, we directly set $\mathcal{I}_{c,r}(t)=\hat{D}_{c,r}(t)$. For top-layer firms with no upstream suppliers, we simply set $\mathcal{I}_{c,r}(t)=\hat{D}_{c,r}(t)$ to focus on resilience inference within the internal supply chain.
7. Inventory balance.
At the end of each time step, inventory levels are updated based on actual inflows and outflows. For finished products, inventory increases with production and decreases with demand fulfillment:
$I_{c,p}(t+1)$ $=I_{c,p}(t)+Q^{\text{plan}}_{c,p}(t)$
$\quad-\min\{I_{c,p}(t),\,D^{\text{agg}}_{c,p}(t)\}$ (13)
For raw materials, inventory increases via upstream inflow $\mathcal{I}_{c,r}(t)$ and decreases through consumption:
$I_{c,r}(t+1)=I_{c,r}(t)+\mathcal{I}_{c,r}(t)-C_{c,r}(t)$ (14)
Resilience Label Generation
For each supply chain instance, we run the simulator under multiple initial inventory conditions. Specifically, we sample $n=10$ random initial states from a log-uniform distribution over $10^{1},10^{3}]$, and add two extreme cases: (i) all inventories set to zero, and (ii) all inventories initialized to a large value (500 units). For each initial condition, we simulate the system for $T=200$ time steps.
The resulting inventory trajectories are compared at the final time step. A binary resilience label $y\in\{0,1\}$ is assigned based on whether all firm–product inventory trajectories converge to a similar steady state. Formally, if the maximum pairwise difference in final inventories is below a threshold $\delta=10$, we assign $y=1$ (resilient); otherwise, $y=0$ (non-resilient).

### v2

@@ Introduction
We represent the supply chain as a tripartite hypergraph $\mathcal{H}=(\mathcal{C},\mathcal{P},\mathcal{E})$, where $\mathcal{C}$ denotes the set of firm nodes, $\mathcal{P}$ is the set of product nodes, and $\mathcal{E}\subseteq\mathcal{C}\times\mathcal{P}\times\mathcal{C}$ represents the set of hyperedges. Each hyperedge $(c_{\text{up}},p,c_{\text{down}})$ indicates that the downstream firm $c_{\text{down}}$ procures product $p$ from the upstream firm $c_{\text{up}}$. Additionally, each firm node $c$ is associated with a time-dependent state vector $\smash{\mathbf{x}_{c}^{(t)}}\in\mathbb{R}^{d}$, capturing its operational status at time $t$. The dimensions of this vector may include key performance indicators such as inventory levels, production rates, and order volumes. In this work, we use inventory vectors to define node states: $\mathbf{x}_{c}^{(t)}=\bigl(I_{c,p}^{(t)}\bigr)_{p\in\mathcal{P}}\in\mathbb{R}^{|\mathcal{P}|}$, where $I_{c,p}^{(t)}$ denotes the inventory level of product $p$ held by firm $c$ at time $t$. Stacking these vectors across all firms forms the supply chain state matrix $\mathbf{X}_{t}\in\mathbb{R}^{|\mathcal{C}|\times|\mathcal{P}|}$.
@@ Acknowledgments
This work was partially supported by the grants of Jilin Provincial International Cooperation Key Laboratory for Super Smart City and Jilin Provincial Key Laboratory of Intelligent Policing, and the Research Institute of Trustworthy Autonomous Systems at Southern University of Science and Technology.